OpenStudy (anonymous):

how do i integrate this 3/x + (x + 1) / (x^2 + 8) u= x^2+8? du=2x?

OpenStudy (anonymous):

Write it as \[ \frac 3 x+ \frac x {x^2+8} + \frac 1 {x^2+8} \] The first two integrals involves natural logairthm in the last one involves \[\tan^{-1} \] For the last integral, I suggest you use substitution with \[ x =\sqrt 8 \tan(u) \]

OpenStudy (anonymous):

hm, so u=sqrt of 8 tan (u) du=

OpenStudy (anonymous):

\[ dx = \sqrt 8 \sec^2(u) du \]